# FedMerge: Federated Personalization via Model Merging

### v2

@@ Abstract
One global model in federated learning (FL) might not be sufficient to serve many clients with non-IID tasks and distributions. While there has been advances in FL to train multiple global models for better personalization, they only provide limited choices to clients so local finetuning is still indispensable. In this paper, we propose a novel “FedMerge” approach that can create a personalized model per client by simply merging multiple global models with automatically optimized and customized weights. In FedMerge, a few global models can serve many non-IID clients, even without further local finetuning. We formulate this problem as a joint optimization of global models and the merging weights for each client. Unlike existing FL approaches where the server broadcasts one or multiple global models to all clients, the server only needs to send a customized, merged model to each client. Moreover, instead of periodically interrupting the local training and re-initializing it to a global model, the merged model aligns better with each client’s task and data distribution, smoothening the local-global gap between consecutive rounds caused by client drift. We evaluate FedMerge on three different non-IID settings applied to different domains with diverse tasks and data types, in which FedMerge consistently outperforms existing FL approaches, including clustering-based and mixture-of-experts (MoE) based methods.
Machine Learning, ICML
1 Introduction
@@
Model Merging (Yadav et al., 2024) is a knowledge ensemble technique that integrates multiple models into a single unified model. It has recently gained popularity in applications involving foundation models (Yang et al., 2024a; Huang et al., 2023; Chronopoulou et al., 2023). Given a large number of foundation models with diverse knowledge domains, model merging facilitates the combination of their knowledge into a new, enriched knowledge space. This approach effectively reuses pre-trained models while avoiding the heavy computational cost of training models from scratch. These reused models can encompass a wide range of techniques, including Parameter-Efficient Fine-Tuning (PEFT) methods such as Low-Rank Adaptation (LoRA). Due to its low computational overhead, LoRA is particularly well-suited for FL, and several studies have explored its applications in this domain (Yi et al., 2023; Yang et al., 2024b; Sun et al., 2024; Cho et al., 2024).
Model merging is especially relevant to FL because it supports both the similarities and differences among the merged models. In FL, clients share commonalities that enable collaborative training while maintaining unique data distributions that reflect their personalized knowledge. This characteristic makes model merging an ideal approach for FL, as it facilitates the integration of shared knowledge across clients by sharing groups of models while preserving individual differences through personalized merging weights.
A number of FL methods with multiple models have been proposed (Yi et al., 2024; Wu et al., 2023; Marfoq et al., 2021a). However, these methods share similar optimization strategies and fail to address the contradiction between the large number of models and the limited resources available to clients. Consider a ”model soup” on the server containing a diverse set of models. Each client utilizes models from this model soup to enhance its personalized training. As shown in Fig. 1(a), previous methods require transmitting the entire model soup to each client, which is often impractical due to the large number of global models. Selecting a subset of models from the model soup (Chronopoulou et al., 2023) can alleviate this issue, but it typically requires task-specific knowledge, making it less generalizable. In contrast, model merging enables each client to leverage the model soup by applying personalized merging weights, thereby forming its own customized model.
@@
Figure 1: Comparison between MoE-like FL methods and the proposed FedMerge. Both method perform weight averaging over global models. MoE-like FL perform weight averaging on each client, while FedMerge perform weight averaging on the server.
In this paper, we propose a novel and general framework for Federated Learning with Model Merging. We introduce a simple merging strategy that decouples the size of the model soup from the constraints of limited client resources. Specifically, instead of transmitting the entire model soup and merging weights to clients, we transmit only a single merged model for each client. As shown in Fig. 1(b), the model merging process is performed on the server rather than on each client, and only the merged models are communicated, eliminating the need to share the original models with clients. In this way, all the computational and storage costs introduced by model merging are shifted to the server, leaving each client to communicate and optimize only a single model. This approach allows us to cheaply increase the size of the model soup on the server to boost performance, while each client only handles a single model.
FedMerge is an end-to-end framework that jointly optimizes the merging weights and global models from scratch. However, FedMerge introduces a unique challenge to FL: model collapse. When the parameters of multiple models are collapsed into a single merged model through weight averaging, clients are optimized only on the merged models, making it impossible to directly update the merging weights or global models on the clients. To address this challenge, we leverage local gradients as intermediate variables and derive a backpropagation-like update strategy for both the merging weights and the global models. This novel update formula forms the basis of the FedMerge algorithm. Experimental results demonstrate that the FedMerge algorithm surpasses traditional FL methods as well as Parameter-Efficient Fine-Tuning (PEFT)-based FL methods on synthetic data distributions and more realistic scenarios.
@@
  • We propose a novel framework for Federated Learning with multiple models, addressing the conflict between large parameter scales and limited client resources.
@@
  • We introduce the Federated Merging (FedMerge) framework, which provides personalized models for each client by merging global models with personalized combining weights.
@@
  • Experimental results demonstrate that FedMerge enables a cost-effective way to improve performance by increasing the number of global models without significantly affecting system costs.
2 Related Work
@@
Model Merging Model merging refers to methods aimed at reusing the knowledge contained in multiple models. The fundamental technique of model merging is to integrate existing models to form new ones. There have been a number of works focusing on merging methodologies, such as simple averaging (Wortsman et al., 2022), merging with Fisher Information (Matena & Raffel, 2022), and merging with task vectors (Ilharco et al., 2022). Recently, model merging has become closely related to the development of foundation models (Sukhbaatar et al., 2024; Lu et al., 2023, 2024; Ostapenko et al., 2024a), as such developments have brought numerous reusable expert foundation models online (Face, 2023). For example, LoraHub (Huang et al., 2023) trains one LoRA expert per task on a collection of 200 tasks and combines these experts to evaluate downstream tasks. Due to the large number of expert models, model selection plays a practical role in choosing the most relevant experts for each downstream task (Zhao et al., 2024; Chronopoulou et al., 2023). Model merging is also related to Mixture of Experts (MoE)-based methods, where the design of routing weights is a key factor. The router can be neural network-based (Lu et al., 2023), feature similarity-based (Wang et al., 2024), or even SVD-decomposed (Ostapenko et al., 2024b). In this paper, rather than directly adopting the various model merging techniques mentioned above in FL, we leverage the fundamental concept of model merging to address the primary challenge in Federated Learning with Multiple Models, enabling a large model soup to be efficiently reused on the server while communicating only a single model to each client.
@@
Federated Learning with Multiple Models The use of multiple models in FL arises from the fact that sharing a single model cannot effectively address the Non-IID problem. Personalized FL methods (Tan et al., 2021; Li et al., 2021; T Dinh et al., 2020; Collins et al., 2021) apply the concept of multiple models by designing a unique model for each client’s personalized data distribution. Our method is more related to a type of multi-model FL, where multiple models are shared on the server (Marfoq et al., 2021a; Wu et al., 2023). In FedEM (Marfoq et al., 2021a), a group of global models is transmitted to each client, and each client optimizes its personalized combination weights. pFedMoE (Yi et al., 2024) employs one local expert and one global expert for each client, while deriving routing weights using personalized router neural networks. Recently, a few methods have incorporated merging concepts into FL (Chen et al., 2024; Salami et al., 2024). However, these methods mainly use model merging as part of their solutions and do not provide a unified framework for model merging in the context of federated learning.
3 Methodology
3.1 Problem Formulation
In Federated Learning, clients collaboratively learn from each other via a central server. Let $(Y_{i},X_{i})$ represent client-$i$’s personalized data. Each client trains local models using $(Y_{i},X_{i})$. Models carrying diverse local knowledge are transmitted from each client to the server, and the server aggregates these local models to centralize the knowledge. The objective of a fundamental FL method, FedAvg, is as follows:
@@ 3.1 Problem Formulation
The $m$ clients’ objective functions are combined based on the proportion of each client’s data $n_{i}$ relative to the total data $n$, collaboratively optimizing a single global model $\Theta$.
In practice, local data across clients are often non-IID (non-identically independently distributed). If we assume that each client owns a personalized model, the personalized FL objective becomes:
@@
where each client optimizes its personalized model $\theta_{i}$ on its own data. The personalized models are often related to some global knowledge.
3.2 Objective Function
In this section, we present the objective and learning algorithm of FedMerge.
The core idea of model merging is to reuse a set of models by combining them to create new ones. Let $\{\Theta_{1},\Theta_{2},\ldots,\Theta_{d}\}$ be a set of $d$ global models, where each model is learned with respect to its own knowledge space. The merging process for creating model $\theta_{i}$ is defined as:
$\theta_{i}=\sum_{j=1}^{d}w_{(i,j)}\cdot\Theta_{j}.$
Here, $w_{i,1:d}$ represents the task-specific merging weights used to create $\theta_{i}$, determining the contribution of each global model to the merged model. By merging global models into new models, their knowledge spaces are reused and integrated, contributing to various downstream tasks.
In federated learning, model merging can be used to address non-IIDness. A model soup containing a set of global models is maintained on the server and reused across all clients. Each client’s personalized model is created by selecting and weight-combining models from the model soup. The objective of FedMerge is given as:
@@ 3.2 Objective Function
[…]
where each client’s personalized model is created by combining a set of global models using its personalized merging weights.
[…]
Figure 2: The update information flow of global models and merging weights for FedMerge. The update information for each global model is derived from the updates of all clients. The update information for each merging weight comes from the global model and merged model it connects.
A commonly employed MoE-like (Mixture of Experts) strategy (Yi et al., 2024; Wu et al., 2023; Marfoq et al., 2021a) in previous methods is to substitute Eq (4) into Eq (3), resulting in directly optimizing $w$ and $\Theta$ on each client. As shown in Fig. 1(a), at each communication round, the MoE-like strategy transmits all the global models to each client, and each client optimizes these global models alongside its personalized merging weights. On the server side, each global model performs FedAvg-like aggregation separately. However, this operation requires broadcasting all the global models to each client, which is impractical when the model soup is large. A potential solution is to use a model selection strategy (Bhuyan & Moharir, 2022), selecting a subset of models from the model soup. While model selection can help alleviate communication costs, it still requires broadcasting multiple models and does not fundamentally solve the problem. The key issue with MoE-like methods is that they omit the merged model $\theta$.
In our method, we assign the merged model $\theta$ the significance it deserves. Specifically, as shown in Fig. 1(b), the merging operation $\theta_{i}=\sum_{j=1}^{d}w_{(i,j)}\cdot\Theta_{j}$ is performed on the server rather than on each client, with both the global models and personalized merging weights stored on the server. Only the merged model $\theta$ is transmitted between the client and the server.
Model merging offers several benefits to federated learning. First, by adjusting the merging weights, it enables the creation of different models tailored to each client’s needs, making it well-suited for heterogeneous federated learning where personalized models are required. Second, by sharing the same group of models, FedMerge allows for a balance between personalization and knowledge sharing, achieving both a certain degree of variation and a certain degree of commonality across clients. Finally, only the merged models are transmitted between the client and the server, which means FedMerge is client-resource friendly. The communicated parameters per round and the optimized parameters per client are equivalent to those in FedAvg. Personalized models are dynamically created on the server, and no local models are remained on each client.
3.3 Optimizing via Backpropagation-like Strategy
Solving Eq (3) with Eq (4) requires optimizing $w$ and $\Theta$. For MoE-like methods, the gradients with respect to $w$ and $\Theta$ are direct and clear: all the global models are sent to each client, and their gradients can be directly computed on each client. However, in the proposed FedMerge, each client only receives the merged model, and only the gradient of the merged model $\frac{\partial\ell}{\partial\theta}$ is directly available. The weights $w$ and global models $\Theta$ remain on the server, and their gradients cannot be computed directly.
The update of merging weights $w$ and global models $\Theta$ is conceptually similar to backpropagation in neural networks. If we consider FedMerge as a single-layer fully connected network, where local models $\theta$ are high-level nodes and global models $\Theta$ are low-level nodes, then $w$ acts as the propagation weights. The gradient of $w$ and $\Theta$ depends on the gradient of $\theta$. Following the chain rule, we compute $\frac{\partial L}{\partial w}$ and $\frac{\partial L}{\partial\Theta}$ using $\frac{\partial L}{\partial\theta}$ as an intermediate variable:
@@ 3.3 Optimizing via Backpropagation-like Strategy
Eq (5) and Eq (6) are derived directly from the objective of FedMerge in Eq (3) with Eq (4). The key factor is the use of local updates $\frac{\partial L}{\partial\theta}$ as an intermediate variable. For the detailed derivation process, please refer to Appendix C. To better illustrate Eq (5) and Eq (6), Fig. 2 highlights the information flows from local models to global models and merging weights. It shows that the updates of each global model are influenced by all client updates, while the updates for each merging weight depend on both the global model and the merged model it connects.
@@
    Forward Propagation: The server computes the merged models for each client using Eq. (4) and distributes them accordingly. Each client then updates its merged model based on its local objective function:
    $\theta^{(t)}_{i}-\theta^{(t-1)}_{i}\leftarrow-\eta_{\theta}\nabla_{\theta^{(t-1)}_{i}}\ell(Y_{i},f(X_{i};\theta^{(t-1)}_{i})),$ (7)
    where $\eta_{\theta}$ is the learning rate for $\theta$. Notably, the local update procedure in FedMerge follows the same approach as in FedAvg.
@@
    Backward Propagation: After the local update, each client transmits its gradient $\frac{\partial\ell}{\partial\theta_{i}}$ or the updated parameters $\Delta\theta_{i}=\theta^{(t)}_{i}-\theta^{(0)}_{i}$ back to the server. The server then updates both the global models $\Theta$ and the merging weights $w$. Using the chain rule, the gradients for the merging weights and global models are computed based on local updates, as formulated in Eq. (5) and Eq. (6).
@@
    Parameter Update: Once the gradients for the global models and merging weights are obtained, the server iteratively updates $\Theta$ and $w$ via gradient descent:
    $\Theta_{j}\leftarrow\Theta_{j}-\eta_{\Theta}\frac{\partial L}{\partial\Theta_{j}},$ (8)
    $w_{(i,j)}\leftarrow w_{(i,j)}-\eta_{w}\frac{\partial L}{\partial w_{(i,j)}},$ (9)
    where $\eta_{\Theta}$ and $\eta_{w}$ are the learning rates for $\Theta$ and $w$, respectively. The overall FedMerge process is summarized in Algorithm 1.
3.4 Technical Implementations
Constraint on Merging Weights Values The original objective of FedMerge in Eq (4) has no constraints on merging weights, allowing the merging weight values to be negative. However, neural network models typically require regularized parameters. To address this, we apply a softmax function to the merging weights to obtain normalized weights. This introduces slight modifications to the gradients in Eq (5) and Eq (6). Detailed derivations can be found in Appendix D.
Clarification of Merging Weights Gradient In Eq (6), the update of merging weights involves the inner product of model parameters and gradients. For large models, the effective update directions can be overwhelmed by the inner products of a large number of parameters. To mitigate this issue, model compression techniques, such as pruning parameters before computing the inner product, can be employed to highlight the most significant gradient directions.
4 Experiments
4.1 Classical FL Settings
Baselines FedMerge is related to Cluster FL, Multi-model FL, and MoE-like FL. To analyze the performance of FedMerge, we consider popular methods as well as state-of-the-art methods as follows: 1) Single-model methods: These FL methods share one model on the server, including Local, FedAvg (McMahan et al., 2017), FedAvg+, and pFedMoE (Yi et al., 2024). Local denotes training local models separately on each client, while FedAvg+ denotes personalized fine-tuning of the global model from FedAvg. Although pFedMoE includes ”MoE” in its name, it uses only one global model. 2) Multi-model methods: These FL methods has flexible number of global models on the server, such as IFCA (Ghosh et al., 2020) and FedEM (Marfoq et al., 2021b). IFCA is a Cluster FL method, and FedEM is a MoE-like FL method.
Initialize : global models $\{\Theta_{1},\Theta_{2},\ldots,\Theta_{d}\}$, merging weight $w\in\mathbb{R}^{m\times d}$
for each communication round do
       Randomly select a subset $A$ of clients;
      Merge models by Eq (4) for each client $i\in A$;
      Send the merged model $\theta_{i}$ to each client-$i$;
      for each selected client in parallel do
            Update $\theta_{i}$ using local gradient descent via Eq (7) for $t$ steps;
            Send local update $\theta^{(t)}_{i}-\theta^{(0)}_{i}$ to the server;
      for each global iteration do
             Compute gradients for $\Theta$ via Eq (5);
            Update $\Theta$ via Eq (8);
            Compute gradients for $w$ via Eq (6);
            Update $w$ via Eq (9);
Algorithm 1 Federated Merging (FedMerge)
Parameter Scale Multi-model FL requires multiple global models. For a fair comparison, we vary the parameter scale across all baselines. For Single-model methods (i.e., Local, FedAvg, FedAvg+, and pFedMoE), we use ResNet-9 (He et al., 2016) as the basic model and vary the number of global models chosen from the set {5, 10, 15, 20, 25, 30}. For Multi-model methods (i.e., IFCA, FedEM, and FedMerge), we vary the model architecture by selecting from ResNet-9, ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152. These models have parameter scales relative to ResNet-9 as follows: {1, 2.27, 4.32, 4.82, 8.73, 11.95}. We summarize our parameter scale settings in Table 1.
Table 1: Parameter scale settings. Model Num. denotes number of models and Model Arch. is model architecture. The scale of parameters is measured as a multiple of the parameter count in ResNet-9.
Multi-Model Model Num. ResNet-9 $\times$ 5 ResNet-9 $\times$ 10 ResNet-9 $\times$ 15 ResNet-9 $\times$ 20 ResNet-9 $\times$ 25 ResNet-9 $\times$ 30

### v3

# FedMerge: Federated Model Merging for Personalization

@@ Abstract
One global model in federated learning (FL) might not be sufficient to serve many clients with non-IID tasks and distributions. Despite recent advances in FL to train multiple global models for better personalization, they only provide limited model choices to clients, so local finetuning of multiple models is still indispensable. This paper proposes a novel “FedMerge” approach that can create a single personalized model per client by simply merging multiple global models with automatically optimized and customized weights. We formulate this problem as a joint optimization of global models and the merging weights per client. Unlike existing FL approaches, where the server broadcasts one or multiple global models to all clients, the server only needs to send a customized, merged model to each client. Moreover, instead of periodically interrupting the local training and re-initializing it to a global model, the merged model aligns better with each client’s task and data distribution, smoothing the local-global gap between consecutive rounds caused by client drift. We evaluate FedMerge on different non-IID settings applied to various domains with diverse tasks and data types, in which FedMerge consistently outperforms existing FL approaches, including clustering-based and mixture-of-experts (MoE) based methods.
Codehttps://github.com/shutong043/FedMerge
Introduction
@@
Figure 1: Comparison between MoE-like FL methods and the proposed FedMerge. Both methods perform weight averaging over global models. MoE-like FL performs weight averaging on each client, while FedMerge performs weight averaging on the server.
Model merging (Yadav et al. 2024) is a knowledge ensemble technique that integrates multiple models into a single unified model. It has recently gained increasing attention in the context of foundation models (Huang et al. 2023; Chronopoulou et al. 2023; Yang et al. 2024a; Kim et al. 2025; Wang et al. 2025), as it enables efficient knowledge integration across diverse domains. Instead of training a large model from scratch, model merging combines existing pre-trained models into an enriched knowledge space, allowing effective reuse of previously trained experts. This flexibility naturally extends to parameter-efficient tuning methods, such as Low-Rank Adaptation (LoRA), which can be merged with minimal computational overhead. Due to these efficiency advantages, LoRA has been widely explored in FL settings (Yi et al. 2023; Yang et al. 2024b; Sun et al. 2024; Cho et al. 2024). These properties make model merging a promising approach for scaling and personalizing federated learning.
Federated learning with multiple models is an important direction for addressing client heterogeneity (Marfoq et al. 2021; Wu et al. 2023), and a recent dominant approach in this line of work is Mixture-of-Experts (MoE) (Luo et al. 2024; Feng et al. 2025; Jiang et al. 2025; Radwan et al. 2025). In these MoE-like FL approaches, the server maintains a “model soup” of diverse experts that are shared across clients. As illustrated in Fig. 1(a), each client downloads multiple (or all) expert models and combines them locally to form a personalized model. However, sharing multiple models across clients introduces an intrinsic conflict: the desire for a rich, diverse set of global models conflicts with the limited computational and communication resources of individual clients. The cost for each client scales with the number of models it downloads, making it impractical to leverage a large model soup. While selecting a subset of models for each client can alleviate this issue (Chronopoulou et al. 2023), it often requires task-specific knowledge, which limits its generalization ability.
Model merging provides an elegant solution to the scalability issues faced by MoE-like FL methods. By integrating multiple expert models into a single personalized model, model merging allows each client to benefit from the richness and diversity of the model soup without having to download or train multiple models locally. This fundamentally eliminates the linear growth of communication and computation costs on the client side, making it feasible to leverage a large number of global models even in resource-constrained environments.
Building on this insight, we propose Federated Learning with Model Merging (FedMerge). As shown in Fig. 1(b), the server merges global models using client-specific weights to construct a unique personalized model for each client. Clients then only need to communicate and optimize a single merged model, while the server handles the merging operation with its typically greater resources. A challenge introduced by this design is how to update both the global models and the merging weights when only merged models are trained on clients. To address this, we formulate FedMerge as a joint optimization problem and derive a backpropagation-like update strategy to enable end-to-end training for both components. Experimental results on various non-IID settings show that FedMerge consistently outperforms clustering-based and MoE-based FL approaches.
@@
  • We propose a novel FL framework to solve the conflict between using multiple global models for personalization and the limited resources of clients.
@@
  • We introduce the Federated Merging (FedMerge) algorithm, where the server provides a personalized model to each client by merging a set of global models with client-specific weights.
  • Convergence analysis is provided in the Appendix E.
@@
  • Experimental results demonstrate that FedMerge offers a cost-effective way to improve performance by supporting multiple global models while maintaining a constant and low client-side cost.
Related Work
@@
Model Merging Model merging refers to methods aimed at reusing the knowledge contained in multiple models. The fundamental technique of model merging is to integrate existing models to form new ones. There have been a number of works focusing on merging methodologies, such as simple averaging (Wortsman et al. 2022), merging with Fisher Information (Matena and Raffel 2022), and merging with task vectors (Ilharco et al. 2022). Recently, model merging has become closely related to the development of foundation models (Lu et al. 2023, 2024; Ostapenko et al. 2024; Zhou et al. 2025; Cheng et al. 2025; Du et al. 2025), as such developments have brought numerous reusable expert foundation models online (Face 2023). For example, LoraHub (Huang et al. 2023) trains one LoRA expert per task on a collection of 200 tasks and combines these experts to evaluate downstream tasks. Due to the large number of expert models, model selection plays a practical role in choosing the most relevant experts for each downstream task (Chronopoulou et al. 2023; Zhao et al. 2024). Model merging is also related to Mixture of Experts (MoE)-based methods, where the design of routing weights is a key factor. The router can be neural network-based (Lu et al. 2023), feature similarity-based (Wang et al. 2024), or even SVD-decomposed (Ostapenko et al. 2024). In this paper, rather than directly adopting the various model merging techniques mentioned above in FL, we leverage the fundamental concept of model merging to address the primary challenge in Federated Learning with Multiple Models, enabling a large model soup to be efficiently reused on the server while communicating only a single model to each client.
@@
Federated Learning with Multiple Models The use of multiple models in FL arises from the fact that sharing a single model cannot effectively address the Non-IID problem. Personalized FL methods (T Dinh et al. 2020; Li et al. 2021; Collins et al. 2021) apply the concept of multiple models by designing a unique model for each client’s personalized data distribution. Our method is more related to a type of multi-model FL, where multiple models are shared on the server (Marfoq et al. 2021; Wu et al. 2023). A recent line of research explores mixture-of-experts (MoE) architectures in FL (Luo et al. 2024; Feng et al. 2025; Jiang et al. 2025; Radwan et al. 2025; Xie et al. 2025), where multiple experts and personalized gating mechanisms are introduced to enhance personalization and handle strong data heterogeneity. These approaches either deploy MoE models on the server to improve scalability or distribute experts across clients for decentralized adaptation, enabling dynamic expert selection and robust personalization across domains. For example, pFedMoE (Yi et al. 2024) employs one local expert and one global expert for each client, while deriving routing weights using personalized router neural networks. Recently, a few methods have incorporated merging concepts into FL (Chen et al. 2024, 2025). In (Tao et al. 2024; Liu et al. 2024), a one-shot federated learning scenario is considered, and model merging methods can be naturally extended to address the challenges rise in this setting. (Salami et al. 2024) proposed an end-to-end optimization framework that leverages RegMean (Jin et al. 2022) to constrain the update distribution of LoRA adapters. While these approaches also utilize model merging to improve federated learning, our work targets a different challenge within this paradigm: balancing the scalability of a large pool of global models with the limited communication and computation capacities of individual clients.
Backgrounds and Motivation
Preliminaries
Federated Learning (FL) enables multiple clients to collaboratively train a model under the coordination of a central server without sharing their private data. Let $(X_{i},Y_{i})$ denote the data of client $i$. In standard FL, the goal is to train a single global model $\Theta$ that minimizes the weighted sum of local objectives across all $m$ clients:
@@
where $n_{i}$ and $n$ denote the data sizes of client $i$ and all clients, respectively. This formulation, known as FedAvg (McMahan et al. 2017), aggregates locally trained models to form a unified global model. However, under non-IID data distributions, a single global model often fails to generalize well to each client’s local domain.
In parallel, the concept of a model soup (Wortsman et al. 2022) refers to a collection of trained models whose parameters are combined—typically via simple averaging or linear interpolation—to improve generalization performance. The core idea is that interpolating between diverse models trained from different optimization trajectories can reveal complementary information, thereby enhancing both robustness and overall generalization.
Problem Setting
To address the limitations of FedAvg under heterogeneous data, personalized federated learning (PFL) allows each client to maintain a customized model $\theta_{i}$, optimized with respect to its local data:
@@
where each $\theta_{i}$ learns client-specific knowledge while still being related to some shared global representation.
Motivated by the insight from model soup, we consider how this notion can be adapted to the PFL setting, where the server can maintain a shared model soup composed of models aggregated or collected from different clients. This shared model soup enables model construction that balance personalization and generalization.
Methodology
Objective Function
We place a model soup containing multiple models on the server. Let $\{\Theta_{1},\Theta_{2},\ldots,\Theta_{d}\}$ denote a set of $d$ global models, where each model is trained with respect to its own knowledge space. In federated learning (FL), each client is associated with a personalized task, and we assume that the knowledge required to solve each task lies within the set of global models maintained on the server, that is, within the shared model soup.
Therefore, the server can create a personalized model for each client by reusing the global models in the model soup. In this work, we adopt standard weight merging, where the global models are combined with task-specific merging weights to generate a personalized model for each client:
$\theta_{i}=\sum_{j=1}^{d}w_{(i,j)}\cdot\Theta_{j},$
where $w_{i,1:d}$ represents the personalized merging weights for client $i$, determining the contribution of each global model to the merged model.
After this merging operation on the server, the personalized merged models are sent to the clients. The objective of FedMerge is formulated as:
@@
$\text{s.t.}\quad\theta_{i}=\sum_{j=1}^{d}w_{(i,j)}\cdot\Theta_{j}.$ (4)
Note that in Eq (4), each client is optimized only on its merged model $\theta_{i}$, rather than directly on the global models or the merging weights. This is a key difference between FedMerge and MoE-like FL methods.
Optimizing via Backpropagation-like Strategy
Solving Eq (3) with Eq (4) requires optimizing $w$ and $\Theta$. In the proposed FedMerge, each client only receives the merged model, and only the gradient of the merged model $\frac{\partial\ell}{\partial\theta}$ is directly available. The weights $w$ and global models $\Theta$ remain on the server, and their gradients cannot be computed directly.
Inspired by the mechanism of backpropagation in neural networks (Rumelhart et al. 1986), we propose an update rule for merging weights $w$ and global models $\Theta$. If we consider FedMerge as a single-layer fully connected network, where local models $\theta$ are high-level nodes and global models $\Theta$ are low-level nodes, then $w$ acts as the propagation weights. The gradient of $w$ and $\Theta$ depends on the gradient of $\theta$. Following the chain rule, we compute $\frac{\partial L}{\partial w}$ and $\frac{\partial L}{\partial\Theta}$ using $\frac{\partial L}{\partial\theta}$ as an intermediate variable:
@@
Eq (5) and Eq (6) are derived directly from the objective of FedMerge in Eq (3) with Eq (4). The key factor is the use of local gradient $\frac{\partial L}{\partial\theta}$ as an intermediate variable. For the detailed derivation process, please refer to Appendix C.
To better illustrate Eq (5) and Eq (6), Fig. 2 highlights the information flows from local models to global models and merging weights. It shows that the gradient of each global model is influenced by all client updates, while the updates for each merging weight depend on both the global model and the merged model it connects.
[…]
Figure 2: The update information flow of global models and merging weights in FedMerge. Global models are updated from all clients, and merging weights are updated from their associated model pairs.
@@
    Forward Propagation: The server computes the merged models for each client using Eq (4) and distributes them accordingly. Each client then updates its merged model based on its local objective function:
    $\theta^{(t)}_{i}-\theta^{(t-1)}_{i}\leftarrow-\eta\nabla_{\theta^{(t-1)}_{i}}\ell(Y_{i},f(X_{i};\theta^{(t-1)}_{i}))$, (7)
    where $\eta$ is the learning rate for $\theta$. Notably, the local update procedure in FedMerge is the same as in FedAvg.
    Algorithm 1 Federated Merging (FedMerge)
    1:Initialize: global models $\{\Theta_{1},\Theta_{2},\ldots,\Theta_{d}\}$, merging weights $w\in\mathbb{R}^{m\times d}$
    2:for each communication round do
    3:  Randomly select a subset $A$ of clients
    4:  for each client $i\in A$ do
    5:   Merge models using Eq (4) to obtain $\theta_{i}$
    6:   Send merged model $\theta_{i}$ to client $i$
    7:  end for
    8:  for all selected clients (in parallel) do
    9:   Update $\theta_{i}$ using local gradient descent via Eq (7) for $t$ steps
    10:   Send local update $\Delta\theta_{i}$ to the server
    11:  end for
    12:  Compute gradients for $\Theta$ via Eq (8)
    13:  Update $\Theta$ via Eq (10)
    14:  Compute gradients for $w$ via Eq (9)
    15:  Update $w$ via Eq (11)
    16:end for
@@
    Backward Propagation: After the local update, each client transmits its updated parameters $\Delta\theta_{i}=\theta^{(t)}_{i}-\theta^{(0)}_{i}$ back to the server. The server then updates both the global models $\Theta$ and the merging weights $w$. Using the chain rule in Eq (5) and Eq (6), the updates for the merging weights and global models are computed based on local updates:
    $\Delta\Theta_{j}=\sum_{i=1}^{m}\frac{n_{i}}{n}w_{(i,j)}\Delta\theta_{i},$ (8)
    $\Delta w_{(i,j)}=\frac{n_{i}}{n}\langle\Theta_{j},\Delta\theta_{i}\rangle.$ (9)
@@
    Parameter Update: Once the gradients for the global models and merging weights are obtained, the server updates $\Theta$ and $w$:
    $\Theta_{j}\leftarrow\Theta_{j}+\Delta\Theta_{j},$ (10)
    $w_{(i,j)}\leftarrow w_{(i,j)}+\Delta w_{(i,j)}.$ (11)
    The overall FedMerge process is summarized in Algorithm 1. We also provide detailed convergence analysis, please refer to Appendix E for details.
Classic FL (ResNet-based)
Multi-Model Model Num. ResNet-9×5 ResNet-9×10 ResNet-9×15 ResNet-9×20 ResNet-9×25 ResNet-9×30